A cube of edge length 10 cm is just balanced at the interface of two liquids A and B . It is 6 cm in liquid B and 4 cm in liquid A. If A and B has specific gravity 0.6 and 0.4 respectively, then the mass of cube is?
(1) 240 g
(2) 360 g
(3) 480 g
(4) 540 g
 
 
Ans. would be option (3)...pls give detailed explanations...

Grade:12th Pass

5 Answers

Khimraj
7 years ago
Balance the force in vertical direction. mg = (pVg)1 + (pVg)2. So m = (0.4 g/ml)600cm³ + (0.6 g/ml)(400 cm³) as 1 ml = 1 cm³ so m = 480 gm.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer
Subhadip Dutta
7 years ago
i’m not getting as to how u r substituting the values...like if u give step by step detailed solutions, then it would be of some help...[ how did u write the vol. as 600 cmand so on...]
Khimraj
7 years ago
As the height of cube emerged in liquid B of specific gravity 0.6 is 6cm so volume will be area of base multiply be height then it is 10*10*6 = 600 cm³. Similarly in liquid A is 400cm³.
Subhadip Dutta
7 years ago
Okk...now i got it...thank u very much..............................................................................................................
ankit singh
3 years ago
 
As the height of cube emerged in liquid B of specific gravity 0.6 is 6cm so volume will be area of base multiply be height then it is 10*10*6 = 600 cm³. Similarly in liquid A is 400cm³.